OpenStudy (nottim):

ADSORPTION OF ACETIC ACID BY ACTIVATED CHARCOAL LAB-"Calculate the exact concentration of your base. To do this, use the volume of base added to 10 mL of your standardized acid solution. Calculate the base concentration for each titration performed then determine the mean concentration."

OpenStudy (nottim):

I'm thinking c1v1=c2v2

OpenStudy (nottim):

I have titration volumes of 11.72, 9.49, 10.27, 10.47, 10.27 mL.

OpenStudy (nottim):

MY Potassium Hydrogen Phthalate concentration is 0.1024 M

OpenStudy (nottim):

1. Freundlich Isotherm 2. Langmuir Isotherm are mentioned in the lab intro.

OpenStudy (dan815):

@hwyl

OpenStudy (nottim):

http://www.isca.in/rjcs/Archives/vol2/i9/9.ISCA-RJCS-2012-145.pdf The proper research article, surprisngly not helpful though.

OpenStudy (anonymous):

is the acid ur titrant or ur base?

OpenStudy (nottim):

i think its the base

OpenStudy (anonymous):

so for ur first one, 11.72 ml of ur base was used to neutralise the acid?

OpenStudy (nottim):

Wait base is the thing that is the beaker already right?

OpenStudy (anonymous):

ive forgotten too haahah, its been a yr since i did titration in uni

OpenStudy (anonymous):

i should know

OpenStudy (anonymous):

but i think you can have either acid or base as ur titrant, it just depends what ur experiment tells u, i think the norm is to have the base in the beaker thought (from my vauge memory)

OpenStudy (nottim):

Alright, any ideas on figuring out concentrations?

OpenStudy (nottim):

test 1

OpenStudy (nottim):

test 2

OpenStudy (anonymous):

best bet is dilution factor c1v1=c2v2

OpenStudy (anonymous):

give me a sec, i'm working on a maths problem as we speak, i posted the question and need some help haha

OpenStudy (nottim):

i dunno what to use for v1 and v2 though i just know that i titrated 11.72 mL into 10mL KHP.

OpenStudy (anonymous):

KHP is ur acid right?

OpenStudy (nottim):

yes

OpenStudy (anonymous):

so you titrated 11.72ml of base into 10ml KHP?

OpenStudy (nottim):

fr my first sample, yes.

OpenStudy (anonymous):

wait, dilution factor is unessecary, do you know the reaction equation?

OpenStudy (nottim):

no, i don't remember what that means

OpenStudy (anonymous):

A+B>C+D etc

OpenStudy (nottim):

oh, no

OpenStudy (nottim):

I do not

OpenStudy (anonymous):

i reckon that the key

OpenStudy (anonymous):

as in do you know the stoichiometry is it 1:1?

OpenStudy (nottim):

I'm looking thru the lab manual, and it says nithing.

OpenStudy (anonymous):

haha well....

OpenStudy (anonymous):

http://www.ausetute.com.au/titrcalc.html

OpenStudy (anonymous):

cause thats sorta what u wanna do, find moles etc then use stoichiometry to find ur base conc

OpenStudy (nottim):

you think i was suppose to deriive that from the equation?

OpenStudy (anonymous):

what is ur base ur using?

OpenStudy (nottim):

sorry, was at bathrooom. its naoh

OpenStudy (anonymous):

well that makes life so much easier ahah

OpenStudy (anonymous):

CH3COOH + NaOH ----> CH3COONa + H2O

OpenStudy (anonymous):

so stoichiometry is 1:1

OpenStudy (anonymous):

C(Acetic)= 0.1024 M V(acetic)=10ml V(naoh)=11.72 C(naoh)=??

OpenStudy (nottim):

wait the acid is khp

OpenStudy (nottim):

acetic is the part after this section of the lab

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[KHC _{8}H _{4}O _{4}+NaOH \rightarrow H _{2}O+NaKC _{8}H _{4}O _{4}\]

OpenStudy (anonymous):

same idea C(KHP)= 0.1024 M VKHP)=10ml V(naoh)=11.72 C(naoh)=??

OpenStudy (anonymous):

C(KHP)= 0.1024 M VKHP)=0.01L V(naoh)=0.01172L C(naoh)=??M

OpenStudy (anonymous):

now, n=CV

OpenStudy (nottim):

lol im stupid i was trying to figure out how you got 0.01172 L...

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

n(kHp)=0.1024*0.01=0.001024 moles

OpenStudy (anonymous):

we know mole ratio is 1:1 therefore number of acid moles=number of base moles

OpenStudy (anonymous):

n(naoh)=0.001024 moles

OpenStudy (anonymous):

n=CV C=n/V C(NaOH)=0.001024/0.01172 C(NaOH)=0.087M

OpenStudy (anonymous):

thats ur first test

OpenStudy (nottim):

i don't know where n=cv is coming from; as in, why its being applied to this scenario

OpenStudy (anonymous):

because u want to find the base concentration...

OpenStudy (anonymous):

and you can only do that using stoichiometry of what u you (acid)

OpenStudy (nottim):

alrihgt.

OpenStudy (anonymous):

so 11.72, 9.49, 10.27, 10.47, 10.27 mL. Test 1: C(naoh)=0.087M Test 2: C(naoh)= 0.001024/(9.49/1000)=0.108M Test 3: C(naoh)= 0.001024/(10.27/1000)=0.0997M Test 4: C(naoh)= 0.001024/(10.47/1000)=0.0978M test 5: C(naoh)= 0.001024/(10.27/1000)=0.0997M Therefore mean concentration is the average right. C(naoh)= test 1+test2+test3+test4+test5/5runs C(naoh)=0.098M (to be honest, no lab supervisor would make up batches of 0.098M, we can safely say that the sodium hydroxide concentration is 0.1M

OpenStudy (nottim):

You just had to explain up to n=CV. But thanks!

OpenStudy (anonymous):

its because moles of a reaction always is constant, its just concentration changes. so we use stoichiometry (which is in effect molar ratio) so we use a comparison of moles to determine unknown concentrations